A man crosses triver perpendicular to river flow in time ‘t’ seconds & travels an equal distance down the stream  in ‘T’ seconds.The ratio of man’s speed in still water to the speed of river will be?

Grade:12th pass

1 Answers

Rishabh singh
6 years ago
Let velocity of man in stillbwater be v, and water speed w.r. t observer in ground be u, So, velocity of man perpendicular moving will be =rootvover v^2-u^2 And velocity of downstream = v+uAcc. To. Que(Rootover v^2-u^2)t= (v+u) T Then v/u comes out to be = t^2+T^2\t^2-T^2